Find the sum of n number of terms.
1/2+3/4+7/8+15/16.......n

Grade:12

1 Answers

Nishant Vora IIT Patna
8 years ago
S= 1/2+3/4+7/8+15/16
S= (1-1/2)+(1-1/4)+(1-1/8)+(1-1/16)+........n terms
S= (1+1+1......n times) - (1/2+1/4+1/8+1/16+........n terms)
S= n - (1/2+1/4+1/8+1/16.....n terms )
Now that is a GP.